# : A Benchmark for Data-Efficiency in Long-Horizon Indoor Mobile Manipulation Robotics

### v6

@@ VI Conclusion
To evaluate the benchmark, we tested multiple baseline models to assess their performance and data efficiency on these complex tasks. End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated extremely poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks. In contrast, a neuro-symbolic system, LIMP, significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by âs tasks. We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in current models and guiding the development of more efficient systems for daily indoor tasks.
Limitations and Future Work. Incorporating a quadruped like Spot in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories. Expanding the benchmark to include simulated quadruped tasks would enhance generalization and enrich evaluation datasets. In the future, the benchmark could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.

### v7

@@ VI Conclusion
To evaluate the benchmark, we tested baseline models to assess their performance and data efficiency on these complex tasks. End-to-end learning-based models, RT-1 and MG-MoMa, demonstrated poor performance and were found to be data-inefficient, highlighting limitations in their ability to generalize to âs long-horizon tasks. In contrast, a neuro-symbolic system, LIMP, significantly outperformed the BC models, showcasing its potential as a promising direction for the development of future MoMa systems capable of addressing the challenges posed by âs tasks. We envision this benchmark as a valuable resource for the robotics community, highlighting data efficiency limitations in models and guiding the development of more efficient MoMa systems.
Limitations and Future Work. Incorporating a quadruped in real-world tasks showcases the advantages of quadrupedal platforms, but the absence of simulated quadruped data limits exposure to diverse environments, such as varied floor-to-floor trajectories. Expanding the benchmark to include that would enhance generalization and enrich evaluation datasets. In the future, could be extended to include tasks beyond pick-and-place, such as those involving more intricate manipulation.